A standard test given to a large number of student had a mean of 74 with a standard deviation of 8.
Find the percentage of students who scored
a. between 72 and 80
b. between 68 and 76
c. more than 76
1
Expert's answer
2011-08-09T10:41:05-0400
a. P(72<X<80) = Ф(3/4) + Ф(1/4) = 0.2734 + 0.0987 = 0.3721
ANSWER:
37.2%

b.& P(68<X<76) = Ф(1/4) + Ф(3/4) = 0.0987 + 0.2734 =
0.3721
ANSWER: 37.2%

c. P(X>76) = 1 - P(X<76) = 1 - Ф(1/4) =
0.9013
ANSWER: 90.1%